OpenStudy (anonymous):

Solve using the quadratic formula: 3x2 = -4x - 5

OpenStudy (anonymous):

discriminant= 16-60=-44 so roots re x=(-4+sqrt(44)i)/6, (-4-sqrt(44)i)/6 sqrt(44)= 2sqrt(11)